1. x+x-4+x-7=19
2. 3x+x-6=54
3. 40=2 (3x+4+x)
HELP

Answer by minerva(22)

1. x+x-4+x-7=19
3x=19+11
x=30/3
x=10
2. 3x+x-6=54
4x=54+6
x=60/4
x=15
3. 40=2 (3x+4+x)
40=2(4x+4)
40=8x+8
40-8=8x
32/8=x
4=x

didnt u just ask word problems, n i answered them??? i thought u just needed the concept.sorry about not being clear